OpenStudy (anonymous):

please heeeelp : suppose f(4)=2, f'(x)=5, g(4)=-3, g'(4)=3. determine each derivative when x=4. a) (fg)(x) b)d/dx (x^2f(x))

OpenStudy (dumbcow):

\[(fg)' = fg' + f'g = 2*3 + 5*-3 = -9\] \[(x^{2} f)' = x^{2}f' + 2x*f = 16*5+8*2 =96\]

OpenStudy (dumbcow):

used the product rule